a. A $1,000 par value bond with a market price of $940 and a coupon interest rate of 7 percent. Flotation costs for a new issue would be approximately 8 percent. The bonds mature in 8 years and the corporate tax rate is 35 percent.

b. A preferred stock selling for $103 with an annual dividend payment of $8.The flotation cost will be $7 per share. The​ company's marginal tax rate is 30 percent.

c. Retained earnings totaling $4.8 million. The price of the common stock is $71 per​share, and dividend per share was $8.24 last year. The dividend is not expected to change in the future.

d. New common stock for which the most recent dividend was $2.83. The​ company's dividends per share should continue to increase at a growth rate of 7 percent into the indefinite future. The market price of the stock is currently $47 ​however, flotation costs of $6 per share are expected if the new stock is issued.
Answer #1
a) Net price realizable after flotation costs = 940-1000*8% = 860
[Flotation rate has been applied on face value]
Coupon rate = 7% (assumed to be annual payment)
n = 8 years
YTM (before tax cost of debt) using an online calculator = 9.58%
After tax cost of debt = 9.58%*(1-35%) = 6.23%
b) Cost of preferred stock = 8/(103-7) = 8.33%
c) Cost of retained earnings = 8.24/71 = 11.61%
d) Cost of new common stock = 2.83*1.07/(47-6)+0.07 = 14.39%